Team:Caltech/Week 7


iGEM 2010



Project Details


Completed Systems



Human Impact



Caltech logo watermark.png

Monday 7/26

CPCR gel featuring L16, L19, L23, L24, L31 & L34. (7/26)
CPCR gel featuring L34, L35, L40, L42, L43 & L44. (7/26)
  • Started PCR reaction to extract the HSP (K112400) from its Berkeley Standard backbone (BBb). (Used same PCR parameters as CPCR from 7/22.
  • Group meeting!
    • By next week: Need to design test-construct experiments, as they will be ready sometime early next week.
      • Can use PCR machine to achieve easy temperature gradients
      • Use spectrophotometer and eventually microscope to assay fluorescence.
      • Consider a plate assay?
    • AIBN issues:
      • Consider using low-concentration DMSO to help dissolve AIBN in LB/cell mixture.
      • Need to figure out how much unreacted (or otherwise) AIBN remains after mixing with cells.
      • Need to test crosslinking using hydrolyzed and epoxidized oil mixed with lysate and with cells to tweak concentration of AIBN needed, and to see what it makes.
    • TGase:
      • purchase urea?
      • To prevent a solid plug forming upon addition of TGase powder to sample, prepare liquid stock of concentrated TGase.
    • Wiki:
      • Should visit for research material for human impact segment.
    • In general, we should set goals for both the end of SURF and for the end of the summer and determine what we need to do to get on track for them.
      • Printing in 2D
      • Working AND gate, at least in bulk fluid
      • Characterization of 3D hydrogel/plastic
      • Crosslinking data
  • PCR purified K112400 PCR product & digested as plasmid insert.
  • Transformed RFP bricks of verious resistances to use as backbone for future ligations
  • Ran 2 gels of today's CPCR products.

Tuesday 7/27

CPCR gel featuring L16, L19, L31, L34, & L35. (7/22)
  • Began more CPCR on failed ones from yesterday: L16, L19, L31, L34, L35
  • Began more digests of miniprepped DNA from K112400
  • Made more electrocompetent cells. Began two test transformations.
  • Transformed ligation products: L42, L46, L47 and the HSP.
  • Began 11 new ligations:
    • L48: R0082 + I13504 + Cm
    • L31: L17 + K215000 + Kan
    • L34: L17 + L21 + Amp
    • L35: L18 + L21 + Amp
    • L22: K156012 + B0034 + Cm
    • L24: K156014 + B0015 + Cm
    • L20: C0077 + B0015 + Cm
    • L49: K112400 + L40 + Cm
    • L25: B0015 + M30109 + Tet
    • L26: R0077 + K124017 + Cm
    • L50: L18 + L27 + Amp
  • Note: Need to upstream-digest more of: J23119, L17, K156014, B0015 for future ligations.

Wednesday 7/28

  • Began CPCR reactions on the HSP, L42, L46, L48.
  • Transformed all ligations from yesterday
  • Re-ran gel from yesterday
    • Same results: looks like 19-4 is a success
  • Ran gel of today's CPCR to determine if the HSP ligations worked.
    • All bands appear to have a 100-400bp smear and no other bands, making the results indeterminate.

Thursday 7/29

  • Re-did the following ligations: L16, L31, L34, L35, HSP (in pSB1C3, for submission), L42, L46, L47
  • Started epic amounts of colony PCR:
    • 3 colonies of each: L42, L46, L31, HSP-Amp, L35, L50, L34, L48, L24, L49, L22, L20
    • Gels for each are included to the right ->
    • Successful: L42-4/-5, L46-4/-5, L47-4, HSP-Amp-5/-6

Friday 7/30

  • Prepped 47-4 (HSP + B0034), L42-4/-5, L46-4/-5 (HSP + R0077), HSP-Amp-5/-6
  • Began a ton of digests:
    • 2x Tet backbone
    • U19: K156012 (us)
    • U20: K156014 (us)
    • U21: C0077 (us)
    • U22: L19 (us)
    • U23: L46 (us)
    • U24: L47 (us)
    • U25: L17 (us)
    • D17: B0034 (ds)
    • D18: B0015 (ds)
    • D19: L18 (ds)
    • D20: L19 (ds)
    • D21: HSP-Amp (ds)

Weekend 7/31-8/1

Saturday 7/31

  • Transformed L22, L24, L33, L43
    • Failed: L51, L52, L53, L25 (cuvette arced)
  • Ran gel of M30109 CPCR product: no bands!
    • Is M30109 incorrect?

Sunday 8/1

  • Removed transformation plates: L22 & L24 were successful; L33 & L53 failed.
Caltech logo.png
Caltech footer.png

Locations of visitors to this page